Shift report 20030518 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun May 18 2003 - 22:56:17 EDT

  • Next message: Ian Bearden: "reducing p+p runs"
    ------------------------------------------------------------
    Date: Sun May 18 22:56:17 2003, Shift: 16:00-24:00
    Supervisor: kh, with: 
    Log book#14 pages 114 - 114
    
    Shift Summary:
    ------------------------------------------------------------
    There was only a few minutes of high background beam during
    the entire shift.  That is a long way to say we got nothing.  
    Then at 22:00 came a call from MCR saying that due to the 
    ongoing vacuum problem with the yellow RF storage cavity that 
    they would "need to go in" (suspecting a broken window) and 
    the estimated time for the fix is 24 HOURS.  So the rest of 
    this shift is cancelled as well as the night shift.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun May 18 2003 - 23:02:15 EDT